The rates of on-time flights for commercial jets are continuously tracked by the U.S. Department of Transportation. Recently, Southwest Air had the best reate with 80 % of its flights arriving on time. A test is conducted by randomly selecting 12 Southwest flights and observing whether they arrive on time. (a) Find the probability that exactly 6 flights arrive late.
Answer #1

Solution

Given that ,

p = 0.80

q = 1 - p = 1 - 0.80=0.20

n = 12

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X = 6) = (12 C 6) * (0.80)6 * (0.20)6

=0.0155

probability=0.0155